Reports 1-1 of 1 Clear search Modify search
AdV-TCS (CO2 laser projector)
fafone, lumaca, pinto, rocchi - 0:05 Tuesday 01 November 2022 (57613) Print this report
NI DAS differential tuning

The goal of the shift was to investigate the possibility to improve the contrast defect with a differential tuning of the DAS.

At the beginning of the shift, there were some difficulties in locking the ITF to CARM null, thus, among the other operations (57614), we have set back the WI DAS to the powers that worked during the weekend: inner @ 73 mW and outer @ 2.09 W at 17:19 UTC.

The ITF reached CARM null soon after and, at the end of the initial transients, we started changing the NI DAS powers @ 18:30 UTC.

The initial configuration is reported in the following table:

  IN [W] OUT [W]
N 0.286 + 0.08 (cold lens) 2.277+1.8 (cold lens)
W 0.073 2.09

The NI DAS powers have been increased by 10% at each step. Here are the details:

  • 18:30 UTC: in from 0.286 W to 0.315 W; out from 2.277 W to 2.505 W;
    • due to a typo, the out DAS was actually increased by 20%, up to 2.675 W;
  • 19:08 UTC: in from 0.315 W to 0.347 W; out unchanged to 2.675 W;
  • 20:15 UTC: in from 0.347 W to 0.382 W; out from 2.675 W to 2.943 W.

We observed that:

  • except for the third step, there is no clear correlation between B1p power and DAS power;
  • the SBs were decreasing since the first step;
  • there is a small increasing trend in the WI DAS power, again connected to the laser head temperature (57612 - to be kept under control).

So, we decided to go back to the initial configuration, and both SBs and B1p improved. Trends are shown in figure 1.

Ongoing simulations show that this behavior is somewhat expected, more details in the next days.

On Wednesday shift, the plan is to tune the WI DAS.

Images attached to this report
Search Help
×

Warning

×